Hi please help,
3/4 PIE + 1/2 COOKIE - 1/4 COOKIE + COOKIE + 1/2 PIE - 1/4 PIE - 3/4 COOKIE.
THANKS
EVELYN

Answer by Mathtut(3670)

3/4 pie+ 1/2 pie -1/4 pie=3/4 pie +2/4 pie -1/4 pie=1 pie
:
1/2 cook -1/4 cook +1 cook -3/4 cook=2/4 cook -1/4 cook +4/4 cook -3/4 cook=2/4=1/2 cookie
:
we converted all fractions to LCD of 4
:
we end up with
:
1 pie + 1/2 cookie